OpenStudy (anonymous):

what does sin(x/2) equal?

OpenStudy (anonymous):

in terms of what?

OpenStudy (anonymous):

in terms of simplifying? eg. sin2x=2sinxcosx. So what is sin(x/2)?

OpenStudy (anonymous):

well then it will be\[\sin \frac{x}{2}=2 \sin \frac{x}{4} \cos \frac{x}{4}\]am i right?

OpenStudy (anonymous):

ye i hope so :P

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so then how would you solve cosx+3sin(x/2) -2 = 0 for 0<=x<=pi?

OpenStudy (anonymous):

i think that will not be useful for this problem, better to write \(\cos x\) in terms of half angles

OpenStudy (anonymous):

so can you explain?

OpenStudy (anonymous):

see, we'r going to make a quadratic out of ur equation

OpenStudy (anonymous):

u sure know the formula\[\cos 2x=\cos^2 x-\sin^2 x\]so\[\cos x=\cos^2 \frac{x}{2}-\sin^2 \frac{x}{2}\]

OpenStudy (anonymous):

so \[\cos x=\color\red{\cos^2 \frac{x}{2}}-\sin^2 \frac{x}{2}=\color\red{1-\sin^2 \frac{x}{2}}-\sin^2 \frac{x}{2}=1-2 \sin^2 \frac{x}{2}\]put this in the equation\[1-2 \sin^2 \frac{x}{2}+3 \sin \frac{x}{2}-2=0\]

OpenStudy (anonymous):

sin pi/2 is 1

OpenStudy (anonymous):

if u let \(u=\sin \frac{x}{2}\) ur equation becomes\[-2u^2+3u-1=0\]which is easy to solve :) makes sense ?

OpenStudy (anonymous):

yes thanks very much :)

OpenStudy (anonymous):

very welcome :)